# arXiv:2602.23189v2[math.AP] 02 Mar 2026

### v2

@@ Abstract
Our method is based on the introduction of suitable modulated quantities and a novel relative entropy functional adapted to the two-phase structure. The key novelty lies in an original comparison of $L^{1}$ and $L^{2}$ norms of the partial densities, exploiting the special structure of the bi-fluid system, together with a variant of Csiszar-Kullback-Pinsker inequality. It allows us to rigorously justify the convergence of weak solutions of the compressible system toward solutions of the corresponding incompressible limit system in the low Mach number regime.
@@ 1 Introduction
Compressible multiphase flow models are a fundamental class of nonlinear systems that arise in fluid mechanics and engineering applications. In many formulations, each phase satisfies its own balance laws, and additional closure relations are required to determine thermodynamic quantities and interfacial effects. Classical two-fluid models developed by Ishii and Hibiki [22], as well as the hyperbolic model introduced by Baer and Nunziato [1], exemplify this framework. Depending on the modeling assumptions, closures may be differential, introducing additional evolution equations, or algebraic, in which the pressure is expressed as a function of partial densities, volume fractions, or other state variables. In particular, algebraic isentropic closures induce a nontrivial coupling between the phases, which has important implications for the analytical structure of the system. From a mathematical perspective, these models have been studied in the compressible PDE community; see, for instance, the handbook chapter by DD. Bresch, B. Desjardins, J.M. Ghidaglia, E. Grenier and M. Hillairet [3], and the references therein. These works provide a mathematical formulation of the models, establish their main structural properties, and contain some of the first analytical results, marking the emergence of sustained interest in such systems within the theory of compressible PDEs.
The growing interest in compressible two-fluid flows with viscosity and weak-regularity data has led to significant mathematical advances, particularly in the viscous framework. Fundamental results on the existence of finite-energy weak solutions were established by Bresch, Mucha, and Zatorska [6] as well as by Vasseur, Wen, and Yu [45], highlighting suitable dissipation mechanisms. In the same spirit, Novotný and Pokorný [39] and Novotný [38] developed a weak solution theory of compressible bi-fluid models, while Jin, Kwon, Nečasová, and Novotný [24] introduced dissipative turbulent solutions and investigated their stability. Weak-strong uniqueness properties were further analyzed by Jin and Novotný [25] and by Li and Zatorska [35]. Moreover, Li and Zatorska [34] addressed weak solutions to the compressible inviscid two-fluid model, emphasizing the additional analytical difficulties arising in the absence of viscosity, whereas Piasecki and Zatorska [41] obtained maximal regularity results in the viscous setting. The large-time behavior of solutions with large initial data was studied by Li, Sun, and Zatorska [33]. Finally, the recent work of Li, Wang, and Zhang [31] established the non-existence of classical solutions in the presence of vacuum, further illustrating the analytical complexity of such models. Note that recently a preprint by Li, Lukáčová-Medvid’ová, Pokornỳ and Zatorska (see [32]) related to weak-strong uniqueness for two-phase systems has been posted on ArXiv independently of ours. These contributions, among others, form a coherent and evolving theory for compressible two-fluid flows.
In this work, we consider a compressible isentropic two-phase model with an algebraic closure and a single common velocity field for both phases. This assumption simplifies the structure of the system while retaining the essential coupling induced by the pressure law. After nondimensionalization, the Mach number $\varepsilon$ appears explicitly in front of the pressure gradient, producing a singular perturbation as $\varepsilon\to 0$. We are interested in the rigorous justification of the corresponding low Mach number limit under well-prepared initial data.
For single-phase flows, the low Mach number limit is by now well understood. For the compressible Euler and Navier–Stokes equations, convergence toward incompressible dynamics under well-prepared initial data has been established in various functional settings; see, for instance, the works of Klainerman and Majda [27], and subsequent developments in the context of symmetric hyperbolic systems and viscous flows.
Without aiming at exhaustiveness, we mention the following representative contributions. In the framework of weak solutions, the low Mach number limit has been justified in [36], [15], [14], and [4]. For strong solutions, one may refer to [21], [37], and [11]. A concise overview of these developments can be found in [19]. A comprehensive treatment of the compressible Navier–Stokes system with constant viscosities, with particular emphasis on the relative entropy method as a stability and convergence tool, is given in [18]. For more complex systems, such as compressible Navier–Stokes equations with density-dependent viscosities, the notion of relative entropy had to be revisited and adapted; see, for instance, [7] and [5]. We also mention works developing the relative entropy framework in related contexts, such as [17] and [43].
In contrast, significantly fewer results are available for multiphase flow models, even in simplified configurations. Only recently has attention been devoted to the formal derivation of low Mach number limits for various two-phase systems, including models with algebraic or PDE closures, as well as formulations involving one or two velocities. In these works, asymptotic expansions are typically performed as the Mach number tends to zero; see, for instance, [44] and [29]. We may also refer to Fanelli et al [16] for related incompressible limits in a reduced compressible MHD system.
In the present paper, in the two-phase setting with equal pressure, we introduce several relevant modulated quantities and propose a relative entropy functional based on a suitable combination of them. This approach requires the development of appropriate comparison tools in order to strengthen certain convergence properties.
Our main result establishes that, in the low Mach limit and under well-prepared initial data, the partial densities converge to constant values, while the velocity field converges to a divergence-free vector field. The volume fractions remain variable and are transported by the limit flow, so the mixture density evolves through advection of the phase fractions. This leads to a variable-density incompressible system in which the mixture density depends only on the transported volume fractions.
1.1 Bi-fluid compressible system with algebraic closure
In the present paper, we analyse a bi-fluid isentropic compressible Navier-Stokes system for barotropic pressures. More precisely, denoting the phases by ++ and -, we are considering the following system.
@@ 1.1 Bi-fluid compressible system with algebraic closure
where $\rho_{i}$, $p_{i}$, $\alpha_{i}$ the density, pressure and volume fraction of phase $i$. The real numbers $\mu$ and $\lambda$ are the constant shear and bulk viscosities verifying, as usual,
@@
Let us notice that, thanks to the algebraic closure, the pressure term in the momentum equation can be simplified:
@@
$u(t,x)\lvert_{\partial\Omega}=0.$ (1.2)
@@
Remark 1.1.
The analysis can also be extended to periodic domains or to two-dimensional settings, possibly under additional assumptions on the initial data.
Let us denote by $\varepsilon$ the Mach number. The incompressible limit is the limit where the speed of the fluid becomes negligible compared to the speed of sound. Thus we reset the flow speed to the scale by replacing $u$ with $\varepsilon u^{\varepsilon}$. A particle will move a distance of order $1/\varepsilon$ during a time of order $1/\varepsilon^{2}$, meaning that the relevant time scale is $1/\varepsilon^{2}$. This is why we can use a change of variable like $t^{\varepsilon}=\varepsilon^{2}t$ and $x^{\varepsilon}=\varepsilon x$, with $t^{\varepsilon}$ and $x^{\varepsilon}$ of order $1$ (see [12]). Thus, we scale $u$ and $\rho$ in the following way
We can rewrite System (LABEL:system) in the following form:
$\left\{\begin{aligned} &\partial_{t}(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon})+\operatorname{div}(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}u^{\varepsilon})=0,\\ &\partial_{t}(\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon})+\operatorname{div}(\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}u^{\varepsilon})=0,\\ &\partial_{t}((\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon})u^{\varepsilon})+\operatorname{div}((\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon})u^{\varepsilon}\otimes u^{\varepsilon})+\boldsymbol{\frac{1}{\varepsilon^{2}}}\nabla p_{+}(\rho_{+}^{\varepsilon})\\ &\hskip 199.16928pt=\mu\Delta u^{\varepsilon}+(\mu+\lambda)\nabla\operatorname{div}u^{\varepsilon},\\ &p_{i}(\rho_{i}^{\varepsilon})=(\rho_{i}^{\varepsilon})^{\gamma_{i}},\qquad\rho_{i}^{\varepsilon}\geqslant 0,\qquad i=+,-\\ &p_{+}(\rho_{+}^{\varepsilon})=p_{-}(\rho_{-}^{\varepsilon}),\\ &\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}=1,\qquad 0\leqslant\alpha_{+}^{\varepsilon}\leqslant 1,\end{aligned}\right$. (1.3)
still considering the Dirichlet boundary condition (1.2).
@@
As usual in the mathematical studies of bi-fluid models, we introduce the fractional masses $R_{\pm}^{\varepsilon}:=\alpha_{\pm}^{\varepsilon}\rho_{\pm}^{\varepsilon}$, the mixture density by $\rho^{\varepsilon}:=R_{+}^{\varepsilon}+R_{-}^{\varepsilon}$ and the momentum by $m^{\varepsilon}:=\rho^{\varepsilon}u^{\varepsilon}$.
Remark 1.2.
@@
For $0\leqslant R_{+}^{\varepsilon}\leqslant\rho_{+}^{\varepsilon}$ and $m^{\varepsilon}$ given, if $\rho_{+}^{\varepsilon}>0$, we may define $\alpha_{+}^{\varepsilon}=R_{+}^{\varepsilon}/\rho_{+}^{\varepsilon}\in[0,1]$. Otherwise, we set $\alpha_{+}=1/2$ on $\{\rho_{+}^{\varepsilon}=0\}$. Next we define $\alpha_{-}^{\varepsilon}=1-\alpha_{+}^{\varepsilon}\in[0,1]$, $\rho_{-}^{\varepsilon}=(\rho_{+}^{\varepsilon})^{\gamma_{-}/\gamma_{+}}\geqslant 0$ (by the equality of pressures), $R_{-}^{\varepsilon}=\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}$ and $\rho^{\varepsilon}=R_{+}^{\varepsilon}+R_{-}^{\varepsilon}$. If $\rho_{+}^{\varepsilon}>0$, we get that $\rho_{-}^{\varepsilon},\rho^{\varepsilon}>0$ and we may define the velocity $u^{\varepsilon}=m^{\varepsilon}/\rho^{\varepsilon}$. Otherwise, we set $u^{\varepsilon}=0$ on $\{\rho_{+}^{\varepsilon}=0\}$.
For any functions $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})\in L^{\gamma_{+}}\times L^{\gamma_{+}}\times L^{1}$ satisfying the following compatibility conditions
$0\leqslant R_{+,0}^{\varepsilon}\leqslant\rho_{+,0}^{\varepsilon}\text{\ a.e.},\quad\sqrt{\rho_{0}^{\varepsilon}}u_{0}^{\varepsilon}\in L^{2}$ (1.4)
@@
where we have defined by the previous remark $\alpha_{\pm,0}^{\varepsilon}\in[0,1]$, $\rho_{-,0}^{\varepsilon}\geqslant 0$, $\rho^{\varepsilon}_{0}\geqslant 0$ and $u_{0}^{\varepsilon}$, we consider a global weak solution “à la Leray”, see Definition 2.1 for the precise meaning.
@@ 1.2 Bi-fluid incompressible system
In the case of nonhomogeneous incompressible flows with constant viscosities, the existence of weak and strong solutions has been investigated in several works. The existence of weak solutions was first established by Kazhikhov [26]. Concerning strong solutions, Ladyzhenskaya and Solonnikov [28] proved local existence, together with global existence for small initial data (and global solvability in two space dimensions). In sufficiently smooth domains, Lemoine [30] obtained local strong regular solutions, which are global for small regular data. Okamoto [40] proved existence and uniqueness of local weak solutions and obtained global solutions in three dimensions under smallness assumptions on the initial data. Simon [42] established the existence of global-in-time weak solutions under suitable weak assumptions on the initial data, and proved global strong solvability provided the initial density is not too small (and sufficiently regular). The solvability of nonstationary problems for nonhomogeneous incompressible fluids, as well as convergence in the vanishing viscosity limit, was studied by Itoh [23]. Choe and Kim [10] established local existence and uniqueness of strong solutions in three dimensions, including situations where the initial density vanishes on an open subset of the domain. More recently, Danchin [13] developed a well-posedness theory in critical Besov spaces and proved the existence of global-in-time unique solutions under suitable smallness assumptions on the initial data.
Among the vast existing literature, let us state a result of local existence of strong solutions of non-homogeneous viscous incompressible Navier-Stokes equations.
@@
Let $(\rho_{0},u_{0})$ be some initial data such that $u_{0}\in H^{1}_{0}(\Omega)\cap H^{s}(\Omega)$, $\operatorname{div}u_{0}=0$ in $\Omega$ and $\rho_{0}\in L^{\infty}(\Omega)$, with $\overline{C}\geqslant\rho_{0}\geqslant\underline{C}>0$ for some constants $\overline{C},\underline{C}$. Then there exists $T>0$ and a (unique) solution $(\rho,u,\Pi)$ of (1.5) such taht $0<\underline{C}\leqslant\rho\leqslant\overline{C}$ that satisfy
$u\in L^{\infty}(0,T;H_{0}^{1}(\Omega)\cap H^{s}(\Omega)),\qquad\frac{\partial u}{\partial t}\in L^{2}(0,T;H_{0}^{1}(\Omega)),$
$\nabla^{s}u\in L^{2}(0,T;L^{6}(\Omega)),\qquad\nabla^{s+1}u\in L^{2}(0,T;L^{3/2}(\Omega)),$
$\rho\in L^{\infty}(0,T;H_{0}^{1}(\Omega)),$
$\frac{\partial\rho}{\partial t}\in L^{\infty}(0,T;L^{2}(\Omega)),\qquad\sqrt{\rho}u\in L^{\infty}(0,T;L^{2}(\Omega)),$
$\nabla\Pi\in L^{\infty}(0,T;L^{2}(\Omega))\cap L^{2}(0,T;L^{6}(\Omega)\cap W^{s-1,3/2}(\Omega)).$
@@
which is well defined for $\gamma_{+}\neq\gamma_{-}$, as $C_{0}>0$. We observe that $\alpha_{\pm}$ satisfy the transport equation. Assuming initially $\alpha_{\pm,0}\in[\underline{a},\bar{a}]\subset(0,1)$, such a property then remains true. Let us notice that we can express $\underline{C}$ and $\overline{C}$ in terms of $\underline{a},\overline{a},C_{0},\gamma_{\pm}$.
@@
We will consider $s=3$. The existence result stated above is not optimal in terms of regularity for our subsequent arguments. For our purposes, a lower level of regularity would suffice, but this is not relevant here.
1.3 Main theorem
We can now state the main result of this paper, where we denote the fractional masses of the limit system by $R_{\pm 0}=\alpha_{\pm,0}\rho_{\pm,0}$.
Theorem 1.6 (Singular limit of the compressible two-phase system).
Let $C>0$, $\gamma_{\pm}\geqslant 2$ be three real numbers and $\alpha_{\pm,0}$ be two functions such that $\gamma_{+}\neq\gamma_{-}$ and $0<\underline{a}\leqslant\alpha_{\pm,0}\leqslant\overline{a}<1$. Let us assume that $(\alpha_{\pm,0},u_{0})$ are enough regular such that we can apply Theorem 1.3 for the initial data $(\rho_{0},u_{0})$, and let $T>0$ be the existence time of strong solution to System (1.5).
For each $\varepsilon>0$, let $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})\in L^{\gamma_{+}}\times L^{\gamma_{+}}\times L^{1}$ such that compatibility conditions (1.4) are verified and the total masses are consistent with the limit:
$\int_{\Omega}R_{\pm,0}^{\varepsilon}\,dx=\int_{\Omega}R_{\pm,0}\,dx>0.$ (1.6)
Assume that
  1. (i)
    the initial energy is uniformly bounded, namely
    $\int_{\Omega}\frac{1}{2}\Big(\rho^{\varepsilon}_{0}|u^{\varepsilon}_{0}|^{2}\Big)(x)\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{+,0}^{\varepsilon}\rho_{+,0}^{\varepsilon\gamma_{+}}(x)}{\varepsilon^{2}(\gamma_{+}-1)}\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{-,0}^{\varepsilon}\rho_{-,0}^{\varepsilon\gamma_{-}}(x)}{\varepsilon^{2}(\gamma_{-}-1)}\,\mathrm{d}x\leqslant C,$ (1.7)
    and
    $\int_{\Omega}R_{+,0}^{\varepsilon}\ln_{+}\frac{R_{+,0}^{\varepsilon}}{R_{+,0}}+\int_{\Omega}R_{-,0}^{\varepsilon}\ln_{+}\frac{R_{-,0}^{\varepsilon}}{R_{-,0}}\leqslant C,$ (1.8)
    where, here and below, $C$ denotes positive constants independent of $\varepsilon$,
  2. (ii)
    the initial data is well prepared, in the sense
    $\begin{split}&\int_{\Omega}\frac{1}{2}\big(R^{\varepsilon}_{+,0}+R^{\varepsilon}_{-,0}\big)|u_{0}^{\varepsilon}-u_{0}|^{2}\,dx\\ &\quad+\int_{\Omega}\frac{\alpha_{+,0}^{\varepsilon}}{\varepsilon^{2}(\gamma_{+}-1)}\Big((\rho_{+,0}^{\varepsilon})^{\gamma_{+}}-\rho_{+,0}^{\gamma_{+}}-\gamma_{+}\rho_{+,0}^{\gamma_{+}-1}(\rho_{+,0}^{\varepsilon}-\rho_{+,0})\Big)\,dx\\ &\quad+\int_{\Omega}\frac{\alpha_{-,0}^{\varepsilon}}{\varepsilon^{2}(\gamma_{-}-1)}\Big((\rho_{-,0}^{\varepsilon})^{\gamma_{-}}-\rho_{-,0}^{\gamma_{-}}-\gamma_{-}\rho_{-,0}^{\gamma_{-}-1}(\rho_{-,0}^{\varepsilon}-\rho_{-,0})\Big)\,dx\longrightarrow 0.\end{split}$ (1.9)
    as $\varepsilon\to 0$.
Then, the corresponding weak solution $(\rho_{\pm}^{\varepsilon},\alpha_{\pm}^{\varepsilon},u^{\varepsilon})$ of the compressible two-phase Navier–Stokes system (LABEL:system_eps) on $(0,T)$, in the sense of Definition 2.1, converge strongly to the incompressible limit as $\varepsilon\to 0$, namely
$R_{\pm}^{\varepsilon}$ $\longrightarrow R_{\pm}$ $\text{strongly in }L^{\infty}(0,T;L^{2}(\Omega)),$ (1.10)
$\rho_{\pm}^{\varepsilon}$ $\longrightarrow\rho_{\pm}$ $\text{strongly in }L^{\infty}(0,T;L^{2}(\Omega)),$ (1.11)
$\alpha_{\pm}^{\varepsilon}$ $\longrightarrow\alpha_{\pm}$ $\text{weak-$*$ in }L^{\infty}(0,T;L^{2}(\Omega)),$ (1.12)
$u^{\varepsilon}$ $\longrightarrow u$ $\text{strongly in }L^{2}(0,T;L^{2}(\Omega)),$ (1.13)
$\nabla u^{\varepsilon}$ $\rightharpoonup\nabla u$ $\text{weakly in }L^{2}(0,T;L^{2}(\Omega)),$ (1.14)
$\operatorname{div}u^{\varepsilon}$ $\rightharpoonup 0$ $\text{weakly in }L^{2}(0,T;L^{2}(\Omega)).$ (1.15)

### v3

# Low-Mach-number limit of a compressible two-phase flow system with algebraic closure

@@ Abstract
Our method relies on the introduction of suitable modulated quantities and on two relative entropy functionals adapted to the two-phase structure: a standard entropy commonly used in the literature, and a logarithmic entropy, which is essential here as the former is not sufficient due to the structure of the underlying two-phase system.
@@ 1 Introduction
This paper addresses the rigorous justification of the low-Mach-number limit for a compressible two-phase flow system. The model under consideration is a compressible isentropic Navier-Stokes system describing the evolution of a mixture of two fluids sharing a common velocity and pressure field, with their densities governed by coupled mass balance equations under barotropic pressure laws.
Physically, the study of two-phase compressible flows in the low Mach number regime is motivated by a wide range of industrial and geophysical applications, where fluids mixtures move at speeds much lower than the speed of sound. Mathematically, the problem provides a natural extension of the classical incompressible limit for single-phase compressible Navier–Stokes equations. However, the presence of several phases introduces additional variables and couplings, resulting in a richer and more complex structure than in the single-phase setting.
In most modeling frameworks for such mixtures, each phase satisfies its own system of balance laws, supplemented by closure relations that account for thermodynamic quantities and interfacial interactions. Classical examples include the two–fluid models developed by Ishii and Hibiki [29] and the hyperbolic system introduced by Baer and Nunziato [3]. Depending on the modeling assumptions, these closures may introduce additional evolution equations or take the form of algebraic relations linking pressure, densities and volume fractions.
The presence of several phases introduces new challenges: the system typically involves several densities and volume fractions transported by the flow, interacting through the pressure and momentum equations. This additional structure makes the rigorous justification of the low-Mach-number limit substantially more delicate, as the coupling between phases complicates the derivation of uniform estimates and the control of nonlinear interactions.
1.1 Main result
For our study we consider a compressible isentropic two-phase flow Navier-Stokes system with algebraic closure. Let us denote by $\varepsilon$ the Mach number. After nondimensionalization, see Section 2.1 below for details, denoting the phases by ++ and -, the system reads as
$\left\{\begin{aligned} &\partial_{t}(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon})+\operatorname{div}(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}u^{\varepsilon})=0,\\ &\partial_{t}(\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon})+\operatorname{div}(\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}u^{\varepsilon})=0,\\ &\partial_{t}((\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon})u^{\varepsilon})+\operatorname{div}((\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon})u^{\varepsilon}\otimes u^{\varepsilon})+\frac{1}{\varepsilon^{2}}\nabla\Big(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+}^{\varepsilon})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-}^{\varepsilon})\Big)\\ &\hskip 199.16928pt=\mu\Delta u^{\varepsilon}+(\mu+\lambda)\nabla\operatorname{div}u^{\varepsilon},\\ &p_{i}(\rho_{i}^{\varepsilon})=(\rho_{i}^{\varepsilon})^{\gamma_{i}},\qquad\rho_{i}^{\varepsilon}\geqslant 0,\qquad i=+,-\\ &p_{+}(\rho_{+}^{\varepsilon})=p_{-}(\rho_{-}^{\varepsilon}),\\ &\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}=1,\qquad 0\leqslant\alpha_{+}^{\varepsilon}\leqslant 1,\end{aligned}\right$. (1.1)
@@ 1.1 Main result
where $\rho_{i}^{\varepsilon}$, $p_{i}$, $\alpha_{i}^{\varepsilon}$ are the density, pressure and volume fraction of phase $i$ and we consider one single velocity $u^{\varepsilon}$. The real numbers $\mu$ and $\lambda$ are the constant shear and bulk viscosities verifying, as usual,
@@
$u^{\varepsilon}(t,x)\lvert_{\partial\Omega}=0$ (1.2)
@@
As usual in the mathematical studies of bi-fluid models, we introduce the fractional masses $R_{\pm}^{\varepsilon}:=\alpha_{\pm}^{\varepsilon}\rho_{\pm}^{\varepsilon}$, the mixture density by $\rho^{\varepsilon}:=R_{+}^{\varepsilon}+R_{-}^{\varepsilon}$ and the momentum by $m^{\varepsilon}:=\rho^{\varepsilon}u^{\varepsilon}$. As emphasized in [9], it is important to be able to pass back and forth between the variables $(\rho_{+}^{\varepsilon},R_{+}^{\varepsilon},m^{\varepsilon})$ and $(\alpha_{\pm}^{\varepsilon},\rho_{\pm}^{\varepsilon},u^{\varepsilon})$.
Definition 1.1.
@@
For $0\leqslant R_{+}^{\varepsilon}\leqslant\rho_{+}^{\varepsilon}$ and $m^{\varepsilon}$ given, if $\rho_{+}^{\varepsilon}>0$, we may define $\alpha_{+}^{\varepsilon}=R_{+}^{\varepsilon}/\rho_{+}^{\varepsilon}\in[0,1]$. Otherwise, we set $\alpha_{+}^{\varepsilon}=1/2$ on $\{\rho_{+}^{\varepsilon}=0\}$. Next we define $\alpha_{-}^{\varepsilon}=1-\alpha_{+}^{\varepsilon}\in[0,1]$, $\rho_{-}^{\varepsilon}=(\rho_{+}^{\varepsilon})^{\gamma_{+}/\gamma_{-}}\geqslant 0$ (by the equality of pressures), $R_{-}^{\varepsilon}=\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}$ and $\rho^{\varepsilon}=R_{+}^{\varepsilon}+R_{-}^{\varepsilon}$. If $\rho_{+}^{\varepsilon}>0$, we get that $\rho_{-}^{\varepsilon},\rho^{\varepsilon}>0$ and we may define the velocity $u^{\varepsilon}=m^{\varepsilon}/\rho^{\varepsilon}$. Otherwise, we set $u^{\varepsilon}=0$ on $\{\rho_{+}^{\varepsilon}=0\}$.
Any functions $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})\in L^{\gamma_{+}}(\Omega)\times L^{\gamma_{+}}(\Omega)\times L^{1}(\Omega)$ are compatibles for the bi-fluid compressible system if
$0\leqslant R_{+,0}^{\varepsilon}\leqslant\rho_{+,0}^{\varepsilon}\text{\ a.e.},\quad\sqrt{\rho_{0}^{\varepsilon}}u_{0}^{\varepsilon}\in L^{2}(\Omega),$ (1.3)
where we have defined $\alpha_{\pm,0}^{\varepsilon}\in[0,1]$, $\rho_{-,0}^{\varepsilon}\geqslant 0$, $\rho^{\varepsilon}_{0}\geqslant 0$ and $u_{0}^{\varepsilon}$ from $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})$.
@@
For such compatible initial data, we consider a global weak solution “à la Leray”, see Definition 2.1 for the precise meaning.
@@
The existence of strong solutions to this system has been extensively studied in the literature (see Section 2.4). For completeness, we recall later an existence result (Theorem 2.3) taken from [52], which specifies the notion of strong solution considered in this work.
In the next theorem, we need to define the notion of well prepared data, which is related to the entropies used. Throughout the article, we will use the following quantities:
  • the energy
    $E^{\varepsilon}(t):=\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u^{\varepsilon}|^{2}\Big)(t,x)\,\mathrm{d}x+\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}(t,x)}{\varepsilon^{2}}H_{i}(\rho_{i}^{\varepsilon}|\rho_{i})(t,x)\,\mathrm{d}x,$ (1.5)
    with, for $i=+,$-,
    $H_{i}(\rho_{i}^{\varepsilon}|\rho_{i}):=\frac{1}{\gamma_{i}-1}\left((\rho_{i}^{\varepsilon})^{\gamma_{i}}-\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\right).$ (1.6)
  • the natural modulated entropy
    ${\mathcal{E}}_{1}^{\varepsilon}(t)$ $={\mathcal{E}}_{1}\Big(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\Big)$ (1.7)
    $=\frac{1}{2}\int_{\Omega}\rho^{\varepsilon}(t,x)|u^{\varepsilon}-u|^{2}(t,x)\,\mathrm{d}x+\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}(t,x)}{\varepsilon^{2}}H_{i}(\rho_{i}^{\varepsilon}|\rho_{i})(t,x)\,\mathrm{d}x.$
  • the Kullback type relative entropy
    ${\mathcal{E}}_{2}^{\varepsilon}(t)={\mathcal{E}}_{2}\Big(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\Big):=\sum_{i=+,-}\int_{\Omega}\Big(R_{i}^{\varepsilon}\ln\frac{R_{i}^{\varepsilon}}{R_{i}}-R_{i}^{\varepsilon}+R_{i}\Big)(t,x)\,\mathrm{d}x.$ (1.8)
Remark 1.2.
Observe that the energy functional $E^{\varepsilon}$ is the natural energy associated with the system. A detailed formal derivation of the corresponding energy identity for smooth solutions can be found in [37], see also Section 2.3. The quantity $\mathcal{E}_{1}^{\varepsilon}$ is also standard (see [44]). The functional ${\mathcal{E}}_{2}^{\varepsilon}$ is less classical, although similar quantities appear in the work of P.-L. Lions on compressible models, where the strict convexity of $x\ln x-x+1$ is used in compactness arguments.
We can now state the main result of this paper, where we denote the fractional masses of the limit system by $R_{\pm}=\alpha_{\pm}\rho_{\pm}$.
@@
[Singular limit of the compressible two-phase system] Let $C_{0}>0$, $\gamma_{\pm}\geqslant 2$ be three real numbers and $\alpha_{\pm,0}$ be two functions such that $\gamma_{+}\neq\gamma_{-}$ and $0<\underline{a}\leqslant\alpha_{\pm,0}\leqslant\overline{a}<1$. Let $(\rho_{+},\rho_{-},\alpha_{+},u)$ be a solution of the non-homogeneous incompressible Navier-Stokes system (1.4), verifying the regularity properties of the limit solution, see Theorem 2.3, and let $T>0$ be the existence time of this solution.
For each $\varepsilon>0$, let $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})\in L^{\gamma_{+}}(\Omega)\times L^{\gamma_{+}}(\Omega)\times L^{1}(\Omega)$ be compatible initial data for the bifluid compressible system in the sense of Definition 1.1, and let $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ be a global weak solution of the compressible two-phase Navier–Stokes system (LABEL:system_eps), in the sense of Definition 2.1.
If we assume that the initial data are well prepared in the sense
${\mathcal{E}}_{1}^{\varepsilon}(0)+{\mathcal{E}}_{2}^{\varepsilon}(0)\to 0\quad\text{as}\quad\varepsilon\to 0\quad\text{and}\quad\Big(\int_{\Omega}R_{\pm,0}^{\varepsilon}\Big)_{\varepsilon}\text{ bounded,}$
then the following holds true when $\varepsilon\to 0$
$\rho_{i}^{\varepsilon}$ $\longrightarrow\rho_{i}$ $\text{strongly in }L^{\infty}(0,T;L^{p}(\Omega)),$
$\alpha^{\varepsilon}_{i}$ $\longrightarrow\alpha_{i}$ $\text{strongly in }L^{q}(0,T;L^{2}(\Omega)),$
$R_{\pm}^{\varepsilon}$ $\longrightarrow R_{\pm}$ $\text{strongly in }L^{q}(0,T;L^{2}(\Omega)),$
$u^{\varepsilon}$ $\longrightarrow u$ $\text{strongly in }L^{2}(0,T;H^{1}_{0}(\Omega)),$
$\rho^{\varepsilon}|u^{\varepsilon}-u|^{2}$ $\longrightarrow 0$ $\text{strongly in }L^{q}(0,T;L^{1}(\Omega)),$
for all $i=+,$-, all $p<\gamma_{i}$ and all $q\in[1,\infty)$.
@@
It will be shown subsequently that the assumption $\int R_{\pm,0}^{\varepsilon},{\mathcal{E}}_{1}^{\varepsilon}(0)$ bounded and ${\mathcal{E}}_{2}^{\varepsilon}(0)\to 0$ yields the convergences $\|R_{\pm,0}^{\varepsilon}-R_{\pm,0}\|_{L^{2}(\Omega)}\to 0$ and $\|\alpha_{\pm,0}^{\varepsilon}-\alpha_{\pm,0}\|_{L^{2}(\Omega)}\to 0$, see Proposition 3.7 and (5.3).
@@
The strong convergences in $L^{\infty}(0,T;L^{p}(\Omega))$ may be quantified by explicit estimates in terms of $\varepsilon$ and ${\mathcal{E}}_{1}^{\varepsilon}(0)$. For the sake of readability, we do not record these bounds in the statement of the theorem, and refer instead to Proposition 3.6 together with the boundedness of ${\mathcal{E}}_{1}^{\varepsilon}$ proved in Section 4.1. In the case where $\gamma_{i}=2$ then the convergence holds for $p\leqslant\gamma_{i}$.